3x-4y+3z=-17
x-3y+2z=-8
x-2y+z=-5
Please help me find all possible solutions (Idenity Matrix)

Answer by Alan3354(69443)

3x-4y+3z=-17
x-3y+2z=-8
x-2y+z=-5
===================================
(-3,-1,-4)